A company's cost function is given as:

TC=25+30Q+0.2Q2 TC = 25 + 30Q + 0.2Q^2 TC=25+30Q+0.2Q2

What is the equation for the company's average variable cost?
Answer #1

Cost function TC=25+30Q+0.2Q^2

TC=TFC+TVC

So,TFC=25

TVC=30Q+0.2Q^2

AVC=TVC/Q=(30Q+0.2Q^2)/Q

=30+0.2Q

Therefore average variable cost=30+0.2Q